# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

### v3

@@ B.3 Pairwise Comparisons
DML-TS-NNR and DML-TS-NNR-BLM perform well across all three settings and the difference between them is statistically indistinguishable in 2 of 3 settings. These methods perform about equally well compared to IntelPooling and NNR-Linear in the first setting, but they perform better in the other two. In the Nonlinear setting in particular, DML-TS-NNR and DML-TS-NNR-BLM substantially outperform all other methods, and the pairwise differences are statistically significant at the 5% level.
@@
1. DML-TS-NNR - 58%* 48% 52% 32%* 100%* 100%* 100%* 70%
2. DML-TS-NNR-BLM 42%* - 40%* 36%* 22%* 100%* 100%* 100%* 63%
3. DML-TS-SU 52% 60%* - 48% 34%* 100%* 100%* 100%* 71%
4. NNR-Linear 48% 64%* 52% - 32%* 100%* 100%* 100%* 71%
5. IntelPooling 68%* 78%* 66%* 68%* - 100%* 100%* 100%* 83%
@@ 8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
1. DML-TS-NNR - 46% 64% 58% 90%* 100%* 100%* 100%* 80%
2. DML-TS-NNR-BLM 54% - 64%* 60% 100%* 100%* 100%* 100%* 83%
3. DML-TS-SU 36% 36%* - 44% 88%* 100%* 100%* 100%* 72%
4. NNR-Linear 42% 40% 56% - 94%* 100%* 100%* 100%* 76%
5. IntelPooling 10%* 0%* 12%* 6%* - 80%* 100%* 100%* 44%
6. Neural-Linear 0%* 0%* 0%* 0%* 20%* - 100%* 100%* 31%
7. Standard 0%* 0%* 0%* 0%* 0%* 0%* - 100%* 14%
@@
1. DML-TS-NNR - 54% 100%* 98%* 98%* 100%* 100%* 100%* 93%
2. DML-TS-NNR-BLM 46% - 94%* 98%* 100%* 100%* 100%* 100%* 91%
3. DML-TS-SU 0%* 6%* - 42% 78%* 100%* 100%* 100%* 61%
4. NNR-Linear 2%* 2%* 58% - 82%* 100%* 100%* 100%* 63%
5. IntelPooling 2%* 0%* 22%* 18%* - 98%* 100%* 100%* 49%
@@
Table 2: Pairwise comparisons between methods in the three settings of the main simulation. Each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. DML-TS-NNR and DML-TS-NNR-BLM perform well in all three settings, and their final regret is statistically indistinguishable in 2 of 3 settings. They substantially outperform all other methods in the Nonlinear setting.
B.4 Simulation with Rectangular Data Array
@@
In this section, we simulate actions and rewards under a rectangular array with 100 users and 100 time points. Although we still follow the staged recruitment regime depicted in Figure 3 in Appendix A, at stage 100 we stop sampling actions and rewards for user 1; at stage 101 we stop sampling for user 2; and so on until we have sampled 100 time points for all 100 users. Aside from the shape of the data array, the setup is the same for this simulation as for the main simulation.
Figure 7: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 3 in Appendix A. Similar to Figure 1, DML-TS-NNR is competitive in the first setting and substantially outperforms the competitors in the other settings.
Figure 8: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 3 in Appendix A. We observe the same performance ordering as in Figure 6, but here the relative differences are even larger.
@@
Figure 8 displays comparisons similar to those shown in Figure 6 but for the rectangular simulation. We again see that the results are qualitatively similar to the main simulation. The methods perform similarly in the Homogeneous Users setting. DML-TS-SU performs poorly in the Heterogeneous Users setting, but the other methods perform similarly. Finally, DML-TS-NNR and DML-TS-NNR-BLM substantially outperform the other methods in the nonlinear setting. Again, the differences in performance are even larger in this simulation than in the main simulation.
@@
1. DML-TS-NNR - 54% 46% 42% 40% 100%* 100%* 100%* 69%
2. DML-TS-NNR-BLM 46% - 46% 38%* 32%* 100%* 100%* 100%* 66%
3. DML-TS-SU 54% 54% - 44% 34% 100%* 100%* 100%* 69%
4. NNR-Linear 58% 62%* 56% - 52% 100%* 100%* 100%* 75%
5. IntelPooling 60% 68%* 66% 48% - 100%* 100%* 100%* 77%
@@ 8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
1. DML-TS-NNR - 56% 100%* 44% 100%* 100%* 100%* 100%* 86%
2. DML-TS-NNR-BLM 44% - 100%* 38% 100%* 100%* 100%* 100%* 83%
3. DML-TS-SU 0%* 0%* - 2%* 92%* 100%* 100%* 100%* 56%
4. NNR-Linear 56% 62% 98%* - 100%* 100%* 100%* 100%* 88%
5. IntelPooling 0%* 0%* 8%* 0%* - 96%* 100%* 100%* 43%
6. Neural-Linear 0%* 0%* 0%* 0%* 4%* - 100%* 100%* 29%
@@
Table 3: Pairwise comparisons between methods in the three settings of the simulation with a rectangular array of data. As in Table 2, each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in term of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. The qualitative results are similar to those of Table 2. DML-TS-NNR and DML-TS-NNR-BLM perform well across all three settings and substantially outperform all other methods in the Nonlinear setting.
Table 3 displays results from pairwise method comparisons, similar to those shown in Table 2. Again the results are qualitatively similar to but slightly more exaggerated than those from the main simulation. In particular, in this simulation study, DML-TS-NNR and DML-TS-NNR-BLM outperform all other methods in 100% of repetitions in the nonlinear setting. One notable difference compared to the main simulation is that NNR-Linear and IntelPooling slightly outperform DML-TS-NNR and DML-TS-NNR-BLM in the first two settings of this simulation.
Appendix C Additional Details for Valentine Study
@@
C.1 Data from the Valentine Study
@@
Since the goal of the Valentine study is to increase participants’ activity levels, we thereby define the reward, $R_{t}$, as the step count for the 60 minutes following a decision point (log-transformed to eliminate skew). Our application also uses a subset of the baseline and contextual data; this subset contains the variables with the strongest association to the reward. Table 4 shows the features available to the bandit in the Valentine study data set.
@@
C.2 Evaluation
The Valentine study collected the sensor-based features at 4 decision points per day for each study participant. The reward for each message was defined to be $\text{log}(0.5+x)$, where $x$ is the step count of the participant in the 60 minutes following the notification. As noted in the introduction, the baseline reward, i.e. the step count of a subject when no message is sent, not only depends on the state in a complex way but is likely dependent on a large number of time-varying observed variables. Both these characteristics (complex, time-varying baseline reward function) suggest using our proposed approach.
We generated 100 bootstrap samples and ran our contextual bandit on them, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. In the case of the algorithm utilizing NNR, we chose four baseline characteristics (gender, age, device, and baseline average daily steps) to establish a measure of “distance" between users. For this analysis, the value of $k$ representing the number of nearest neighbors was set to 5. To utilize bootstrap sampling, we train the Neural-Linear method’s neural network using out-of-bag samples. The neural network architecture comprises a single hidden layer with two hidden nodes. The input contains both the baseline characteristics and the contextual variables and the activation function applied here is the softplus function, defined as $\text{softplus}(x)=\log(1+\exp{(x)})$.
@@ C.2 Evaluation
C.3 An ablation study
We further investigate the primary contributors to the algorithm’s performance, we conducted a parallel ablation study using real-world data analysis. In this study, we compared DML-TS-NNR (as detailed in the main paper) with two additional methods for reference: DML-TS-SU and NNR-Linear. Each comparison focuses on specific elements of our algorithm to discern their impact on DML-TS-NNR’s overall performance.
@@ C.3 An ablation study
C.4 Inverse Propensity Score (IPS) offline evaluation
@@
$\hat{R}_{\text{IPS}}=\frac{1}{T}\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}r_{t}$ (11)
@@
To address the instability issue caused by re-weighting in some cases, we use a Self-Normalized Inverse Propensity Score (SNIPS) estimator. This estimator scales the results by the empirical mean of the importance weights, and still maintains the property of unbiasedness.
@@
Appendix D Additional Details for the Intern Health Study (IHS)
To further enhance the competitive performance of our proposed DML-TS-NNR algorithm, we conducted an additional comparative analysis using a real-world dataset from the Intern Health Study (IHS) (NeCamp et al., 2020). This micro-randomized trial investigated the use of mHealth interventions aimed at improving the behavior and mental health of individuals in stressful work environments. The estimates obtained represent the improvement in average reward relative to the original constant randomization, averaging across stages (K = 30) and participants (N = 1553). The available IHS data consist of 20 multiple-imputed data sets. We apply the algorithms to each imputed data set and perform a comparative analysis of the competing algorithms. The results presented in Figure 10 shows our proposed DML-TS-NNR algorithm achieved significantly higher rewards than the other three competing ones and demonstrated comparable performance to the AC algorithm. These findings further support the advantages of our proposed algorithm.
D.1 Data from the IHS

### v4

@@ A.3 Pairwise Comparisons
RoME and RoME-BLM perform well across all three settings and the difference between them is statistically indistinguishable. These methods perform about equally well compared to IntelPooling and NNR-Linear in the first setting, but they perform better in the other two. In the Nonlinear setting in particular, RoME and RoME-BLM substantially outperform all other methods, and the pairwise differences are statistically significant at the 5% level.
@@
1. RoME - 58%* 46% 66% 40% 100%* 100%* 100%* 73%
2. RoME-BLM 42%* - 36%* 44% 32%* 100%* 100%* 100%* 65%
3. RoME-SU 54% 64%* - 58% 44% 100%* 100%* 100%* 74%
4. NNR-Linear 34% 56% 42% - 32%* 100%* 100%* 100%* 66%
5. IntelPooling 60% 68%* 56% 68%* - 100%* 100%* 100%* 79%
@@ 8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
1. RoME - 58% 80%* 56% 100%* 100%* 100%* 100%* 85%
2. RoME-BLM 42% - 72%* 52% 100%* 100%* 100%* 100%* 81%
3. RoME-SU 20%* 28%* - 22%* 92%* 100%* 100%* 100%* 66%
4. NNR-Linear 44% 48% 78%* - 100%* 100%* 100%* 100%* 81%
5. IntelPooling 0%* 0%* 8%* 0%* - 94%* 100%* 100%* 43%
6. Neural-Linear 0%* 0%* 0%* 0%* 6%* - 98%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 2%* - 100%* 15%
@@
1. RoME - 56% 100%* 96%* 98%* 100%* 100%* 100%* 93%
2. RoME-BLM 44% - 96%* 98%* 100%* 100%* 100%* 100%* 91%
3. RoME-SU 0%* 4%* - 40% 80%* 100%* 100%* 100%* 61%
4. NNR-Linear 4%* 2%* 60% - 82%* 100%* 100%* 100%* 64%
5. IntelPooling 2%* 0%* 20%* 18%* - 98%* 100%* 100%* 48%
@@
Table 3: Pairwise comparisons between methods in the three settings of the main simulation. Each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in terms of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. RoME and RoME-BLM perform well in all three settings, and their final regret is statistically indistinguishable. They substantially outperform all other methods in the Nonlinear setting.
A.4 Simulation with Rectangular Data Array
@@
In this section, we simulate actions and rewards under a rectangular array with 100 users and 100 time points. Although we still follow the staged recruitment regime depicted in Figure 1, at stage 100 we stop sampling actions and rewards for user 1; at stage 101 we stop sampling for user 2; and so on until we have sampled 100 time points for all 100 users. Aside from the shape of the data array, the setup is the same for this simulation as for the main simulation.
Figure 7: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 1. Similar to Figure 2, RoME is competitive in the first setting and substantially outperforms the competitors in the other settings.
Figure 8: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings using a rectangular array of data in which we observe 100 time points for 100 users in a stagewise fashion as depicted in Figure 1. We observe the same performance ordering as in Figure 6, but here the relative differences are even larger.
@@
Figure 8 displays comparisons similar to those shown in Figure 6 but for the rectangular simulation. We again see that the results are qualitatively similar to the main simulation. The methods perform similarly in the Homogeneous Users setting. RoME-SU performs poorly in the Heterogeneous Users setting, but the other methods perform similarly. Finally, RoME and RoME-BLM substantially outperform the other methods in the nonlinear setting. Again, the differences in performance are even larger in this simulation than in the main simulation.
@@
1. RoME - 58% 44% 54% 48% 100%* 100%* 100%* 72%
2. RoME-BLM 42% - 42% 48% 40% 100%* 100%* 100%* 67%
3. RoME-SU 56% 58% - 60% 54% 100%* 100%* 100%* 75%
4. NNR-Linear 46% 52% 40% - 52% 100%* 100%* 100%* 70%
5. IntelPooling 52% 60% 46% 48% - 100%* 100%* 100%* 72%
@@ 8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
1. RoME - 52% 100%* 52% 100%* 100%* 100%* 100%* 86%
2. RoME-BLM 48% - 100%* 52% 100%* 100%* 100%* 100%* 86%
3. RoME-SU 0%* 0%* - 2%* 92%* 100%* 100%* 100%* 56%
4. NNR-Linear 48% 48% 98%* - 100%* 100%* 100%* 100%* 85%
5. IntelPooling 0%* 0%* 8%* 0%* - 88%* 100%* 100%* 42%
6. Neural-Linear 0%* 0%* 0%* 0%* 12%* - 100%* 100%* 30%
@@
Table 4: Pairwise comparisons between methods in the three settings of the simulation with a rectangular array of data. As in Table 3, each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in terms of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret. The qualitative results are similar to those of Table 3. RoME and RoME-BLM perform well across all three settings and substantially outperform all other methods in the Nonlinear setting.
Table 4 displays results from pairwise method comparisons, similar to those shown in Table 3. Again, the results are qualitatively similar to but slightly more exaggerated than those from the main simulation. In particular, in this simulation study, RoME and RoME-BLM outperform all other methods in 100% of repetitions in the Nonlinear setting.
A.5 Hyperparameter Sensitivity
In the simulation study, we chose hyperparameters for the fairest comparison possible, employing the same regularization parameters across methods. To assess robustness, we performed seven sensitivity analyses that alter hyperparameters for our methods while fixing those of the other methods; this approach gives the competing algorithms an advantage. Under these alternative hyperparameters, RoME and RoME-BLM still substantially outperform the other methods in the Nonlinear Setting.
Two of the sensitivity analyses led to no significant changes to Figure 2 and Table 3: setting $\delta=0.05$ and $v=10$. Below, we summarize the changes in the remaining analyses:
  • Rescaling $\gamma$ by a factor of ten: This change resulted in RoME and RoME-BLM performing relatively better compared to the other methods in the Heterogeneous setting. In particular, RoME and RoME-BLM outperformed NNR-Linear in 70% and 72% of replications, compared to 56% and 52% using the original value of $\gamma$.
  • Rescaling $\lambda$ by a factor of ten: Similar to the above results, this changed improved the performance of RoME and RoME-BLM compared to the other methods (especially RoME-SU and NNR-Linear) in the Heterogeneous Users setting.
  • Adding (low and medium) noise to the nearest-neighbor network: RoME significantly outperformed RoME-BLM in the Nonlinear setting.
  • Increasing the number of neighbors to 10: The performance gap between the full RoME algorithms (RoME and RoME-BLM) and RoME-SU increased to 88%, 76% compared to 80% and 72% under the original configuration, presumably due to increased network cohesion.
The results for the $\lambda$ sensitivity analysis are displayed in Table 5. Results for the remaining sensitivity analyses are available with our code.
Homogeneous Users
1 2 3 4 5 6 7 8 Avg
1. RoME - 58% 48% 64% 38% 100%* 100%* 100%* 73%
2. RoME-BLM 42% - 38%* 42% 38%* 100%* 100%* 100%* 66%
3. RoME-SU 52% 62%* - 58% 46% 100%* 100%* 100%* 74%
4. NNR-Linear 36% 58% 42% - 32%* 100%* 100%* 100%* 67%
5. IntelPooling 62% 62%* 54% 68%* - 100%* 100%* 100%* 78%
6. Neural-Linear 0%* 0%* 0%* 0%* 0%* - 98%* 100%* 28%
7. Standard 0%* 0%* 0%* 0%* 0%* 2%* - 100%* 15%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Heterogeneous Users
1 2 3 4 5 6 7 8 Avg
1. RoME - 52% 92%* 72%* 100%* 100%* 100%* 100%* 88%
2. RoME-BLM 48% - 78%* 68%* 100%* 100%* 100%* 100%* 85%
3. RoME-SU 8%* 22%* - 20%* 92%* 100%* 100%* 100%* 63%
4. NNR-Linear 28%* 32%* 80%* - 100%* 100%* 100%* 100%* 77%
5. IntelPooling 0%* 0%* 8%* 0%* - 96%* 100%* 100%* 43%
6. Neural-Linear 0%* 0%* 0%* 0%* 4%* - 96%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 4%* - 100%* 15%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Nonlinear
1 2 3 4 5 6 7 8 Avg
1. RoME - 52% 96%* 96%* 98%* 100%* 100%* 100%* 92%
2. RoME-BLM 48% - 92%* 96%* 96%* 98%* 100%* 100%* 90%
3. RoME-SU 4%* 8%* - 60% 90%* 100%* 100%* 100%* 66%
4. NNR-Linear 4%* 4%* 40% - 82%* 100%* 100%* 100%* 61%
5. IntelPooling 2%* 4%* 10%* 18%* - 98%* 100%* 100%* 47%
6. Neural-Linear 0%* 2%* 0%* 0%* 2%* - 98%* 100%* 29%
7. Standard 0%* 0%* 0%* 0%* 0%* 2%* - 100%* 15%
8. AC 0%* 0%* 0%* 0%* 0%* 0%* 0%* - 0%
Table 5: Pairwise comparisons between methods in a sensitivity analysis in which we scale $\lambda$ by a factor of ten. As in Table 3, each cell indicates the percent of repetitions (out of 50) in which the method listed in the row outperformed the method listed in the column in terms of final regret. Asterisks indicate p-values below 0.05 from paired two-sided t-tests on the differences in final regret.
Appendix B Additional Details for Valentine Study
@@
B.1 Data from the Valentine Study
@@
Since the goal of the Valentine study is to increase participants’ activity levels, we define the reward, $R_{t}$, as the step count for the 60 minutes following a decision point (log-transformed to eliminate skew). Our application also uses a subset of the baseline and contextual data; this subset contains the variables with the strongest association with the reward. Table 6 shows the features available to the bandit in the Valentine study data set.
@@
B.2 Justification of Assumption 4
The motivation for Assumption 4 arises from an exploratory analysis we performed using data from the Valentine study. We constructed the pseudo-reward for each observation as suggested by Equation (1) and then conducted an ANOVA test. The results show clear heterogeneity between users and across time, motivating the adoption of user- and time-specific random effects. The results of the ANOVA test are displayed in Table 7. Figure 9 shows the shape of the heterogeneity in the treatment effects over time.
Df Sum Sq Mean Sq F value Pr(>F)
ParticipantIdentifier 107 264 2.464 8.09 <2e-16 ***
Week 25 17 0.683 2.24 4e-04 ***
Residuals 2265 690 0.305
Table 7: ANOVA analysis of the pseudo-outcomes in the Valentine study. The small p-values constitute strong evidence that the treatment effects differ by participant and over time.
Figure 9: The time heterogeneity in the pseudo-outcomes. We calculated the pseudo-outcomes using (1), then averaged them across participants and plotted them over time. This exploratory analysis shows evidence that the causal effects vary substantially over time.
B.3 Evaluation
The Valentine study collected the sensor-based features at 4 decision points per day for each study participant. The reward for each message was defined to be $\text{log}(0.5+x)$, where $x$ is the step count of the participant in the 60 minutes following the notification. As noted in the introduction, the baseline reward (the expected step count of a subject when no message is sent), not only depends on the state in a complex way but is likely dependent on a large number of time-varying observed variables. Both of these characteristics (complex, time-varying baseline reward function) suggest using our proposed approach.
We generated 100 bootstrap samples and ran our contextual bandit on them, considering the binary action of whether or not to send a message at a given decision point based on the contextual variables $S_{1}$ and $S_{2}$. Each user is considered independently and with a cohesion network, for maximum personalization and independence of results. To guarantee that messages have a positive probability of being sent, we only sample the observations with notification randomization probability between $0.01$ and $0.99$. In the case of the algorithm utilizing NNR, we chose four baseline characteristics (gender, age, device, and baseline average daily steps) to establish a measure of “distance” between users. For this analysis, the value of $k$ representing the number of nearest neighbors was set to 5. To utilize bootstrap sampling, we train the Neural-Linear method’s neural network using out-of-bag samples. The neural network architecture comprises a single hidden layer with two hidden nodes. The input contains both the baseline characteristics and the contextual variables and the activation function applied here is the softplus function, defined as $\text{softplus}(x)=\log(1+\exp{(x)})$.
@@ B.3 Evaluation
B.4 An ablation study
We further investigate the primary contributors to the algorithm’s performance, we conducted a parallel ablation study using real-world data analysis. In this study, we compared RoME (as detailed in the main paper) with two additional methods for reference: RoME-SU and NNR-Linear. Each comparison focuses on specific elements of our algorithm to discern their impact on RoME’s overall performance.
@@ B.4 An ablation study
B.5 Inverse Propensity Score (IPS) offline evaluation
@@
$\hat{R}_{\text{IPS}}=\frac{1}{T}\sum_{t=1}^{T}\frac{\pi_{t}(a_{t}|s_{t})}{p_{t}(a_{t}|s_{t})}r_{t}$
@@
To address the instability issue caused by reweighting in some cases, we use a Self-Normalized Inverse Propensity Score (SNIPS) estimator. This estimator scales the results by the empirical mean of the importance weights, and still maintains the property of unbiasedness.
@@
Appendix C Additional Details for the Intern Health Study (IHS)
To further enhance the competitive performance of our proposed RoME algorithm, we performed an additional comparative analysis using a real-world data set from the Intern Health Study (IHS) (NeCamp et al., 2020). This micro-randomized trial investigated the use of mHealth interventions aimed at improving the behavior and mental health of individuals in stressful work environments. The estimates obtained represent the improvement in average reward relative to the original constant randomization, averaging across stages (K = 30) and participants (N = 1553). The available IHS data consist of 20 multiple-imputed data sets. We apply the algorithms to each imputed data set and perform a comparative analysis of the competing algorithms. The results presented in Figure 11 shows our proposed RoME algorithm achieved significantly higher rewards than the other three competing ones and demonstrated performance comparable to the AC algorithm. These findings further support the advantages of our proposed algorithm.
C.1 Data from the IHS